OpenStudy (anonymous):

i have 6(7+4) and i got 66 but i have to rewrite each expression. what does it mean thanks

OpenStudy (kirbykirby):

I can't think of anything else other than = 6(11) = 66 or = 6(7) + 6(4) = 42 + 24 = 66 I'm not sure what else it could mean

OpenStudy (anonymous):

thanks you

OpenStudy (anonymous):

can i add you as a fan

OpenStudy (kirbykirby):

Sure :)

OpenStudy (anonymous):

thank you for the medal how do you give medals

OpenStudy (kirbykirby):

Just as you did now ;)

OpenStudy (anonymous):

oh okay lol